6.) A 11.5 ml solution of HCl has 0.531 g/L of HCl in solution. Calculate the concentration, amount of moles of HCl, and pH of the solution.
1
Expert's answer
2020-05-07T14:09:08-0400

1)C(HCl)=0.531/36.5=0.0145 mol/L

2)n(HCl)=0.531*0.0115/36.5=1.673 *10^-4 mol

3)pH=-lg[H+]=-lg(1.673*10^-4)=3.78